### v1

# Squeezing and Entanglement Dynamics in Phase-Sensitive Non-Hermitian Systems

Abstract
Over the past decade, parity-time (PT) symmetry and anti-PT (APT) symmetry in various physical systems have been extensively studied, leading to significant experimental and theoretical advancements. However, physical systems that simultaneously exhibit both PT and APT symmetry have not yet been explored. In this study, we construct a phase-sensitive non-Hermitian wave mixing model that inherently possesses single-mode APT symmetry. By tuning the phase of the pump field, this model simultaneously exhibits two-mode quadrature-PT symmetry. Moreover, the APT phase transition is accompanied by the emergence of quantum entanglement from absence to presence. This remarkable quantum effect, quantum entanglement, which distinguishes itself from classical physics, is surprisingly linked to APT symmetry phase transition. Our work further explores the relationship between two-mode quantum entanglement and the phase of the pump field, offering deeper insight into the generation and evolution of entanglement in the corresponding nonlinear system. This provides a new perspective on quantum information processing.
preprint: APS/123-QED
Introduction.—Although any Hermitian operator has real eigenvalues, hermiticity is not a necessary condition to obtain real eigenvalues. Recent studies have shown that any Hamiltonian satisfying parity-time (PT) symmetry ($[H,PT]=0$) can still possess real eigenvalues [1, 2, 3]. Such systems undergo a phase transition during which the PT symmetry is spontaneously broken. The eigenvalues of the Hamiltonian transition from real to complex at a specific point in the parameter space are known as the exceptional point (EP). At the EP, both the eigenvalues and eigenvectors of the Hamiltonian coalesce, leading to a range of novel phenomena [4, 5, 6, 7, 8, 9, 10]. Corresponding to PT symmetry, anti-PT (APT) symmetry, as another form of non-Hermitian symmetry ($\{H,PT\}=0$), also undergoes spontaneous APT symmetry breaking during the APT phase transition, accompanied by the emergence of EPs [11, 12].
Over the past decade, PT symmetry [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and APT symmetry [26, 27, 28, 29, 30, 31, 32, 33] in various physical systems such as photonics, acoustics, and ultracold atoms have been extensively explored, leading to significant experimental and theoretical advancements. To realize PT symmetry in a physical system, the system typically exhibits linear gain and loss [3, 4, 26, 34]. However, these optical gain and loss mechanisms present significant challenges for achieving PT symmetry in the purely quantum realm [35] because the associated Langevin noise disrupts PT symmetry in quantum systems [36, 37]. Currently, there are two approaches to circumvent this issue: one is a passive scheme based on loss modulation, and the other is the Hamiltonian dilation method [37, 38, 39, 40, 41, 42]. By contrast, APT symmetry can be realized without gain and loss [29, 43, 44, 45]. Recent theoretical research indicates that APT symmetry can be realized in a two-mode bosonic system without involving linear gain and loss of classical fields. This suggests the potential for achieving quantum APT symmetry free from Langevin noise, which can be implemented in nonlinear wave-mixing processes [45].
Recent studies focusing on quantum entanglement [46, 47] have demonstrated the unique properties of PT symmetry and APT symmetry in both two-qubit systems [48] and linear optical systems [49]. However, physical systems that simultaneously exhibit both PT symmetry and APT symmetry have not yet been explored.
In this study, we construct a phase-sensitive non-Hermitian wave mixing model that not only possesses inherent single-mode APT symmetry but also achieves two-mode quadrature-PT symmetry by modulating the phase of the pump field. This aspect has not been explored in previous studies. Classical-nonclassical polarity (CNP) reveals the unique impact of APT symmetry on the nonclassical properties of a system. Specifically, significant changes in the single-mode squeezing dynamics and two-body Gaussian entanglement dynamics of the system are observed as it traverses EP. Furthermore, the results from the inseparability criterion provide deeper insights into the relationship between two-mode quadrature-entanglement and the phase of the pump field. They reveal the influence of single-mode APT symmetry and two-mode quadrature-PT symmetry on the entanglement. Specifically, when the phase of the pump field satisfies $\cos\phi=0$, the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry. Under these conditions, the two-mode quadrature-entanglement undergoes sudden death in the two-mode quadrature-PT-broken region and EP, while exhibiting periodic revival characteristics in the two-mode quadrature-PT-symmetric region. Conversely, when the phase of the pump field meets the condition $\cos\phi=\pm 1$, the phase transition of the system from the APT-symmetric region to the APT-broken region is accompanied by the emergence of quantum entanglement from nonexistence (symbolizing a transition from classical to quantum). Our work further investigates the unique non-Hermitian properties of nonlinear systems, exploring the relationship between two-mode quantum entanglement and the phase of the pump field. This deeper examination reveals the entanglement generation and evolution process in the corresponding nonlinear system of the model, offering new perspectives for quantum information processing.
Theoretical model.—Consider the Four-wave mixing process shown in Fig. 1(a). The input pump field $\omega_{p}$ with phase $\varphi$ and coherent state input signal field $\omega_{s}$ (idler field $\omega_{i}$) are incident onto a nonlinear medium (where $\chi$ represents the nonlinear susceptibility of the medium). During this process, two pump photons are annihilated and one signal photon and one idler photon are simultaneously generated. Considering an undepleted pump field, in the classical parametric approximation, the coupling equations between the signal field operator $a_{s}$ and idler field operator $a_{i}$ can be written as (we set $\hbar=1$)[12, 29]
$i\partial_{z}(a_{s},a_{i}^{\dagger})^{T}=H_{\mathrm{single-mode}}(a_{s},a_{i}^{\dagger})^{T}.$ (1)
The non-Hermitian Hamiltonian matrix $H_{\mathrm{single-mode}}$ in Eq. (1) is
$H_{\mathrm{single-mode}}=\left(\begin{array}[]{cc}\Delta&\kappa e^{-i\phi}\\ -\kappa e^{i\phi}&-\Delta\end{array}\right)$ (2)
where $\phi=2\varphi$, $\kappa$ is the real nonlinear coupling coefficient, and $\mathrm{\Delta}=-\Delta k/2$. As shown in Fig. 1(a), the phase mismatch $\Delta k=2k_{p}-\left(k_{s}+k_{i}\right)\cos\theta$, where $k_{p}$, $k_{s}$, and $k_{i}$ represent the wave numbers of the pump $(\omega_{p})$, signal $(\omega_{s})$, and idler fields $(\omega_{i})$, respectively. Obviously, no matter the value of the phase $\phi$, $H_{single-mode}$ satisfies $\left\{H_{single-mode},PT\right\}=0$, where $P$ is the parity operator and $T$ is the time-reversal operator. This indicates that the single-mode system exhibits APT symmetry. Since there is no involvement of loss in $H_{single-mode}$, the Langevin noise operator does not need to be considered here to preserve the commutation relations $[a_{j}(0),a_{j}^{\dagger}(0)]=[a_{j}(L),a_{j}^{\dagger}(L)]=1(j=s,i)$. The eigenvalues of $H_{single-mode}$ are given by $\lambda_{\pm}=\pm\lambda=\pm i\left|\kappa\right|\sqrt{1-\delta^{2}}$, where $\delta=|\mathrm{\Delta}/\kappa|$. As shown in Fig. 1(b), when $\delta<1$, the eigenvalues are a pair of purely imaginary complex conjugates and the system is in the APT-symmetric region. When $\delta>1$, the eigenvalues are real and the system is in the APT-broken region. When $\delta=1$, the system is at the EP, where $\lambda_{0}=0$, indicating spontaneous symmetry breaking in the system.
Figure 1: (a) A non-Hermitian system composed of a four-wave mixing process and a phase-controllable pump source. $\chi$ represents the nonlinear susceptibility. $\varphi$ is the phase of the pump field, $\theta$ is the angle between the pump field and the signal (idler) field, and $L$ represents the total length of the nonlinear medium along the optical axis of the pump field. (b) APT phase transition process of the eigenvalues $\lambda_{\pm}$.
Interestingly, by combining the quadrature of the optical field $q_{j}=\frac{1}{2}(a^{\dagger}_{j}+a_{j})$ and $p_{j}=\frac{i}{2}(a^{\dagger}_{j}-a_{j})$, where $\left[q_{j},p_{j}\right]=\frac{i}{2}$, the dynamical equation between the two-mode quadrature can be rewritten as
$i\partial_{z}(q_{s}+q_{i},p_{s}+p_{i})^{T}=H_{\mathrm{two-mode(+)}}(q_{s}+q_{i},p_{s}+p_{i})^{T},$ (3a)
$i\partial_{z}(p_{s}-p_{i},q_{s}-q_{i})^{T}=H_{\mathrm{two-mode(-)}}(p_{s}-p_{i},q_{s}-q_{i})^{T}.$ (3b)
In here, if and only if $\cos\phi=0$, the corresponding effective Hamiltonians in Eqs. (3a) and (3b) are given by
$H_{two-mode(+)}=\left(\begin{array}[]{cc}-i\kappa\sin\phi&i\Delta\\ -i\Delta&i\kappa\sin\phi\end{array}\right)$ (4c)
$H_{two-mode(-)}=\left(\begin{array}[]{cc}-i\kappa\sin\phi&-i\Delta\\ i\Delta&i\kappa\sin\phi\end{array}\right)$ (4f)
which satisfying $\left[H_{two-mode(\pm)},PT\right]=0$, the two-mode effective Hamiltonian exhibits PT symmetry. The corresponding eigenvalues are given by $\lambda_{two-mode(\pm)}=\lambda_{\pm}=\pm\left|\kappa\right|\sqrt{\delta^{2}-1}$. When $\delta>1$, the eigenvalues are real and the system is in the two-mode quadrature-PT-symmetric region. When $\delta<1$, the eigenvalues are a pair of purely imaginary complex conjugates and the system is in the two-mode quadrature-PT-broken region. When $\delta=1$, the system is at the EP, where $\lambda_{0}=0$, indicating spontaneous symmetry breaking. The solutions to Eqs. (1) and (3a)-(3b) can be found in the supplementary materials.
It is worth emphasizing that by satisfying a specific phase condition ($\cos\phi=0$), this phase-sensitive non-Hermitian wave mixing model can simultaneously exhibit single-mode APT symmetry and two-mode quadrature-PT symmetry, which has not been explored in previous research.
Classical-Nonclassical Polarity.—To investigate the intrinsic connection between non-classical properties and the non-Hermitian nature of this phase-sensitive model, we first study Classical-Nonclassical Polarity (CNP) [50] in our model.
Using the covariance matrix $\gamma_{si}$ (see supplementary materials for more detail), the single-mode and two-mode CNPs can be written as [50]
$P_{1}=-\frac{\kappa^{4}\sin^{4}\lambda L}{\lambda^{4}},$ (5a)
$P_{2}=-\frac{\left(\kappa^{2}-2\Delta^{2}+\kappa^{2}\cos 2\lambda L\right)\kappa^{2}\sin^{2}\lambda L}{\lambda^{4}}.$ (5b)
$P_{1}$ characterizes the single-mode squeezing property of the system. In our model, the signal and idler fields exhibit the same single-mode squeezing properties. $P_{2}$ characterizes the two-mode Gaussian entanglement of the system. When the CNP is greater than 0, it indicates that the system is in a nonclassical state. When the CNP is equal to zero, the system is at the boundary between the classical and nonclassical states. When the CNP is less than 0, the system is in a classical state. The absolute value of the CNP quantifies the distance of the state from the classical-nonclassical boundary.
The nonclassical properties measured by CNP have two important characteristics. First, as shown in Eq. (5a) and (5b), the intensity of the input coherent state is independent of the nonclassical properties measured by the CNP. Second, the phase of the pump field is also independent of the nonclassical properties measured by the CNP, meaning that single-mode squeezing and two-mode Gaussian entanglement are not affected by the phase of the pump field. The nonclassical properties primarily depend on the system parameters $\kappa$ and $\mathrm{\Delta}$, as well as the evolution distance $\kappa L$. Therefore, discussing the CNP connection between single-mode APT symmetry and two-mode quadrature-PT symmetry here does not violate the condition $\cos\phi=0$.
Figure 2: CNP of the system. The evolution of $P_{1}$ (a) and $P_{2}$ (b). The red solid line (for $\delta=0.8$) is in the APT-symmetric region, the blue solid line (for $\delta=1$) is at the EP, and the orange solid line (for $\delta=1.2$)) is in the APT-broken region.
Fig. 2 (a) and (b) describe the evolution of $P_{1}$ and $P_{2}$, respectively. In fig. 2(a), regardless of the single-mode APT symmetry state, $P_{1}$ is less than or equal to 00, indicating that the system does not exhibit single-mode squeezing, and thus demonstrates classical characteristics. Correspondingly, regardless of how the two-mode quadrature-PT symmetry varies, $P_{2}$ is greater than or equal to 00 in fig. 2(b), indicating that the system consistently generates two-mode Gaussian entanglement throughout and reflecting nonclassical characteristics.More deeply, in single-mode APT-symmetric region and two-mode quadrature-PT-broken region ($\delta=0.8$), one of the eigenmodes disappears after a long evolution length owing to the presence of purely imaginary eigenvalues $\lambda_{\pm}$. The two-mode Gaussian entanglement measured by $P_{2}$ shows a monotonic increase. In the single-mode APT-broken region and two-mode quadrature-PT-symmetric region($\delta=1.2$), the two-mode Gaussian entanglement exhibits oscillatory behavior with a period of $T=\kappa\pi/\lambda$. At the EP ($\delta=1$), the two eigenmodes merge into a single mode. The two-mode Gaussian entanglement increases monotonically, but the rate of enhancement is slower compared to the single-mode APT-symmetric region and two-mode quadrature-PT-broken region. CNP reveals the unique connection between single-mode APT symmetry and two-mode quadrature-PT symmetry in the system’s non-classical properties.
The inseparability criterion.—To further explore the system’s nonclassical properties, we consider another characterization of quantum entanglement: the inseparability criterion [51, 52]. It reveals the relationship between system entanglement and the phase of the pump field and elucidates the impact of single-mode APT symmetry and two-mode quadrature-PT symmetry on entanglement.
The inseparability criterion is defined as follows: $E_{1}=Var\left(q_{s}-q_{i}\right)+Var\left(p_{s}+p_{i}\right)$, $E_{2}=Var\left(q_{s}+q_{i}\right)+Var\left(p_{s}-p_{i}\right)$. $E_{1}$ and $E_{2}$ are referred to as type-I and type-II entanglements of the system, respectively. $E_{1}\ <1$ or $E_{2}<1$ corresponds to the weaker criterion (wc). When $E_{1}\ <0.5$ or $E_{2}<0.5$, it corresponds to the stronger criterion (sc). The closer the values of $E_{1}$ and $E_{2}$ are to 0, the higher the degree of entanglement.
The theoretical calculation results are as follows:
$E_{1}=\cos^{2}\lambda L+\frac{\left(\kappa^{2}+\Delta^{2}+G1\right)\sin^{2}\lambda L+G2}{\lambda^{2}}$ (6a)
$E_{2}=\cos^{2}\lambda L+\frac{\left(\kappa^{2}+\Delta^{2}-G1\right)\sin^{2}\lambda L-G2}{\lambda^{2}}$ (6b)
Where $G1=2\kappa\Delta\cos\phi$ and $G2=\kappa\lambda\sin 2\lambda z\sin\phi$. Consider the input field as coherent state. From Eqs. (6a) and (6b), it is evident that the degree of entanglement is independent of the intensity of the input field. Instead, it is determined by the system parameters $\kappa$ and $\mathrm{\Delta}$ and the evolution distance $\kappa L$, and is also influenced by the phase of the pump field. Under different phase conditions (e.g., $\cos\phi=\pm 1$ or $\cos\phi=0$), the system exhibits distinct entanglement evolution characteristics.
Figure 3: The evolution of the type-I/type-II entanglement under different phase $\phi$: ($a_{1}$), ($b_{1}$), and ($c_{1}$) correspond to type-I entanglement $E_{1}$; ($a_{2}$), ($b_{2}$), and ($c_{2}$) correspond to type-II entanglement $E_{2}$. ($a_{1}$) and ($a_{2}$) show the system in the APT-symmetric region with $\delta=0.8$; ($b_{1}$) and ($b_{2}$) show the system at the EP with $\delta=1$; ($c_{1}$) and ($c_{2}$) show the system in the APT-broken region with $\delta=1.2$. The dashed gray line indicates that the evolution simultaneously satisfies two-mode quadrature-PT symmetry($\cos\phi=0$).
Fig.3 shows the evolution of type-I/type-II entanglement in different phase $\phi$. When the phase $\phi$ satisfies $\cos\phi=0$ (e.g., $\phi=1.5\pi$ or $0.5\pi$), the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry, as shown by the dashed gray line. By comparing fig.3 ($\mathrm{a}_{1}$), ($b_{1}$), and ($c_{1}$) with fig.3 ($a_{2}$), ($b_{2}$), and ($c_{2}$), it can be seen that the type-I entanglement $E_{1}$ and type-II entanglement $E_{2}$ differ only by a phase of $\pi$. Therefore, the following analysis focuses on the type-I entanglement $E_{1}$. Fig.3($a_{1}$) shows that when the system is in the single-mode APT-symmetric region ($\delta=0.8$) and the phase $\phi$ satisfies $\pi<\phi<2\pi$, the system is still capable of generating entanglement. However, as the system evolves, entanglement is maintained within a very narrow phase range. Outside this range, entanglement briefly exists before rapidly vanishing. When the phase $\phi$ assumes other values (e.g., $0<\phi<\pi$), no entanglement is generated in the system. Fig.3($c_{1}$) shows that in the single-mode APT-broken region ($\delta=1.2$), the evolution of the system entanglement exhibits an oscillatory pattern with a period $T=\kappa\pi/\lambda$. The degree of entanglement can be tuned by varying the phase $\phi$. Interestingly, as shown in Fig.3($b_{1}$), when the system is at the EP ($\delta=1$), by tuning the phase $\phi$ close to 0, entanglement can be maintained over a wider phase range than in the single-mode APT-symmetric region. This demonstrates the significant role of EP control in non-Hermitian systems and phase $\phi$ tuning in the generation of entanglement in this model.
Specifically, when the system is at the EP ($\delta=1$), the expressions for $E_{1}$ and $E_{2}$ are simplified to
$E_{1}=1+2\kappa^{2}L^{2}-2\kappa^{2}L^{2}\cos\phi+2\kappa L\sin\phi$ (7a)
$E_{2}=1+2\kappa^{2}L^{2}+2\kappa^{2}L^{2}\cos\phi-2\kappa L\sin\phi$ (7b)
It can be observed that when the phase condition $\cos\phi=1$ (or $\cos\phi=-1$) is met, $E_{1}=1$ (or $E_{2}=1$) is independent of the propagation distance, indicating that the system is in a critical state between the classical and quantum regimes.
Figure 4: The evolution of entanglement when the pump phase satisfies $\cos\phi=1$. The solid red line represents the single-mode APT-symmetric region ($\delta=0.8$), the solid blue line indicates the EP ($\delta=1$), the solid orange line denotes the single-mode APT-broken region ($\delta=1.2$), the solid black line represents the weaker criterion (wc) for entanglement, and the dashed black line corresponds to the stronger criterion (sc) for entanglement. (a) Evolution of type-I entanglement. (b) Evolution of type-II entanglement.
Fig. 4 clearly reveals the unique phenomenon that occurs in the system when the phase of the pump field satisfies the condition $\cos\phi=1$. As shown in Fig. 4(a), for type-I entanglement $E_{1}$, at the EP ($\delta=1$), $E_{1}=1$ represents a critical state between the classical and quantum regimes. No entanglement was observed in the single-mode APT-symmetric region ($\delta=0.8$). In the single-mode APT-broken region ($\delta=1.2$), the system consistently exhibits and maintains a certain degree of entanglement, with the entanglement degree oscillating with a period $T=\kappa\pi/\lambda$. The phase transition of the system from the APT-symmetric region through the EP to the APT-broken region is accompanied by a transition of quantum entanglement from nonexistence to presence, symbolizing a shift from classical to quantum behavior. The remarkable quantum effect of quantum entanglement, which distinguishes it from classical physics, is linked to APT symmetry phase transition. For type-II entanglement as shown in fig. 4(b), no entanglement is generated under this phase setting.
Conclusion.—In summary, unlike previous studies, this study introduces a phase-sensitive non-Hermitian four-wave mixing model. The single-mode effective Hamiltonian of this model satisfies APT symmetry, and by adjusting the phase of the pump field, the two-mode effective Hamiltonian can simultaneously exhibit two-mode quadrature-PT symmetry. By analyzing the system’s non-classical characteristics through CNP and the inseparability criterion, it is found that the single-mode squeezing and two-mode Gaussian entanglement measured by CNP are independent of the phase of the pump field, whereas the two-mode quadrature-entanglement measured by the inseparability criterion is phase-dependent. This is primarily because CNP measures the overall non-classicality of the system, whereas the inseparability criterion assesses the entanglement between quadrature components of the two-mode fields. Consequently, they reveal different aspects of non-classical properties.
Theoretical calculations of the inseparability criterion show that, when the pump phase satisfies $\cos\phi=0$, the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry. Under these conditions, quantum entanglement undergoes sudden death in the two-mode quadrature-PT-broken region and EP, whereas it exhibits periodic revival characteristics in the two-mode quadrature-PT-symmetric region. When the pump phase satisfies the condition $\cos\phi=\pm 1$, the phase transition of the system from the APT-symmetric region through the EP to the APT-broken region is accompanied by a transition of quantum entanglement from nonexistence to existence, symbolizing a shift from classical to quantum behavior. Our work further explores the unique non-Hermitian properties of nonlinear systems, investigates the relationship between two-mode quantum entanglement and the pump phase, and provides a deeper understanding of the entanglement generation and evolution processes in the model. This study offers new insights into quantum information processing.
Acknowledgments.—This work was supported by the Innovation Program for Quantum Science and Technology (Grant No.2021ZD0303401), the Quantum Science Strategic Project of Guangdong Province (Grant No.GDZX2306004), the National Natural Science Foundation of China under (No.Grant 62375089 and No.Grant 62471188), and Guangdong Basic and Applied Basic Research Foundation (Grant No.2022A1515140139).